OpenStudy (anonymous):

it takes 5 hrs for 6 cooks to prepare food , how long will it take 7 cooks to prepare food?

OpenStudy (anonymous):

this is a proportion. 5/6 = x/7 x= 35/6

OpenStudy (anonymous):

How much is 35/6 in hours ?

OpenStudy (anonymous):

5.83333333333 Hrs

OpenStudy (anonymous):

or 5 hours and 50 min

OpenStudy (anonymous):

Thanks

hero (hero):

So it's going to take more time for seven cooks?

hero (hero):

It only took five hours for six cooks, but it's going to take almost 6 hours for seven cooks? How does that make sense?

hero (hero):

@GregTheo

OpenStudy (anonymous):

oh yeah... misread the question

OpenStudy (anonymous):

$5400 is invested. at 11% and 7%. whats 11% of 5400? whats 7% of 5400?

hero (hero):

@elo099, what about the question about the cooks?

OpenStudy (anonymous):

oh, um ... idk

OpenStudy (anonymous):

5 hrs at 6 cooks is 30 cook hours. if 7 cooks work then 30/7 is the number of hours

OpenStudy (anonymous):

OK So this can be a little difficult to understand -- so let's show it this way: If 6 cooks do it on 5 hours, how long would it take 1 cook? 5 hours = time for one cook / 6 cooks 30 = time for one cook So if it takes 1 cook 30 hours, how long would it take 8 cooks: 30 hours for 1 cook / 8 cooks = 3.75 hours or 3 hours and 45 minutes. Hope this explanation is helpful.

OpenStudy (anonymous):

I think its inverse variation, since if you have more cooks it takes less time. 5 = k/6 k = 30 x = 30/7 = 4.28

hero (hero):

^ What @niravshah08 said

OpenStudy (anonymous):

@KNicksfan4life you copy pasted it wrong :P

OpenStudy (anonymous):

So r all of you guys hella smart in math???

OpenStudy (anonymous):

you answer will be 4 hrs 17 mins

OpenStudy (anonymous):

Please post new questions separately

OpenStudy (anonymous):

594

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Yes, Thanks